Suppose that 500 tickets were sold for a luncheon, each adult ticket costs 2.50 and each child cost 1.50, if the cash reciet totaled 880. how many tickets were sold of each?
Answer by srinivas123(23)

Suppose that 500 tickets were sold for a luncheon, each adult ticket costs 2.50 and each child cost 1.50, if the cash reciet totaled 880. how many tickets were sold of each?
let the number of adult tickets sold be 'x'.
Then the child tickets sold will be '500-x'.
SO ,
2.5(x) + 1.50(500-x) = 880.
==> 2.5x-1.5x = 880-750.
==> x=130.
So adult tickets = 130 . and child tickets = 370.